# Specialization after Generalization: Towards Understanding Test-Time Training in Foundation Models

### v2

@@ 1 Introduction
Since the “ImageNet moment” in 2012 when AlexNet won the ImageNet challenge (Krizhevsky et al., 2012), scaling data, parameters, and compute have led to foundation models that achieve impressive performance on a wide range of tasks. This has spurred research on scaling laws, suggesting that scaling pre-training of a single model on a broad data distribution is sufficient for good performance on downstream tasks (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022). With first-generation foundation models, fine-tuning was used primarily to adapt models to out-of-distribution test data (i.e., with a distribution shift) or to leverage fresh training data that was not seen during pre-training (so-called “privileged” data). Test-time training (TTT; Sun et al., 2020; Hardt & Sun, 2024; Akyürek et al., 2025) emerged as pushing this mechanism to the extreme: fine-tuning a separate model for each prediction. In recent years, foundation models have grown so large that most test data is effectively “in-distribution”, meaning the model has encountered similar data during pre-training. This raises a key question:
@@
Our work posits that today’s foundation models are “underparameterized” (Kaplan et al., 2020; Bubeck & Sellke, 2021), as evidenced by the continuing improvements in performance when scaling models. We hypothesize that due to this underparameterization, even if test data is in-distribution, the model cannot simultaneously approximate the ground truth across the full data distribution. TTT offers a mechanism to specialize the model to a local area around the test example. By temporarily “forgetting” irrelevant pre-trained knowledge, the model “frees up” capacity to learn the relevant concepts to the immediate task at a higher resolution. We refer to this mechanism as specialization after generalization. The mechanism of TTT—temporarily reallocating capacity by “forgetting” irrelevant knowledge—connects to concepts of capacity saturation and interference studied in continual learning (McCloskey & Cohen, 1989; Kirkpatrick et al., 2017).
@@
The LRH can be formalized as follows. We assume the existence of an $s$-sparse concept space $\Phi:\mathcal{X}\to\mathbb{R}^{d_{1}}$} which is approximated by a learned feature map $\Psi:\mathcal{X}\to\mathbb{R}^{d_{2}}$} with $d_{2}\ll d_{1}$. The ground truth to be learned is linear in the concept space, i.e., $f(x)=\langle\Phi(x),w_{\star}\rangle$ for some unknown $w_{\star}\in\mathbb{R}^{d_{1}}$. [^1] In classification, the logits are canonically parameterized as $\langle\Phi(x),w_{c}\rangle$ where $w_{c}$ is the class-specific weight vector. The class probability is then given by $\mathbb{P}(c\mid x)\propto\exp(\langle\Phi(x),w_{c}\rangle)$. In Section˜3, we leverage the LRH to develop a mechanistic understanding of how TTT behaves, and make the following key observations:
@@ 2. O2:
    Among a test point $x^{\star}\in\mathcal{X}$ and its neighborhood (in $\Psi$-space), $f$ can be approximated by an $s$-sparse linear function in the concept space $\Phi$.
@@ 2 Related work
In the classical machine learning paradigm, models are trained on a fixed training set and then kept frozen during evaluation. Despite this standard practice that was used for decades, early work suggested specializing the model at test-time to each prediction task—such examples are local learning (Cleveland, 1979; Cleveland & Devlin, 1988; Atkeson et al., 1997) and local fine-tuning (Bottou & Vapnik, 1992). More recently, the idea of TTT (Sun et al., 2020; Wang et al., 2021) has regained attention in the context of fine-tuning large foundation models during evaluation (e.g., Krause et al., 2018; Hardt & Sun, 2024; Sun et al., 2024). TTT for a few gradient steps on (self-)supervised losses has since shown success in domains such as control (Hansen et al., 2021), abstract reasoning (Akyürek et al., 2025; Zweiger et al., 2025), language modeling (Hardt & Sun, 2024; Hübotter et al., 2025; Sun et al., 2024; von Oswald et al., 2025; Yu et al., 2025), and video generation (Dalal et al., 2025). Many standard TTT methods train on carefully selected data from the pre-training dataset (i.e., do not add any new privileged information; Hardt & Sun, 2024; Hübotter et al., 2025), and several works studied how to optimally select data for imitation, e.g., the early seminal work of MacKay (1992) and recent extensions (Hübotter et al., 2024; Bagatella et al., 2025b). TTT has also been extended from supervised learning to reinforcement learning (Zuo et al., 2025; Bagatella et al., 2025a; Diaz-Bone et al., 2025).
@@
Our theoretical framework is built upon the LRH, which posits that foundation models represent high-level concepts as linear directions in their activation spaces. This idea has its roots in early word embedding models, which famously showed that semantic analogies could be solved with simple vector arithmetic (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016). More recently, the LRH has been validated across a wide range of models and domains, with studies identifying linear representations for abstract concepts like sentiment (Tigges et al., 2023), the state of a game board (Nanda et al., 2023), and even fundamental axes of space and time (Gurnee & Tegmark, 2024). A key tool for discovering and studying these conceptual directions are SAEs (Makhzani & Frey, 2014; Lieberum et al., 2024; Gao et al., 2025). SAEs are auxiliary models trained to reconstruct a foundation model’s internal activations from a sparse, overcomplete dictionary of features. This process often yields features that are monosemantic, or aligned with single, human-interpretable concepts, thereby providing an empirical method to uncover the sparse concept space we consider in our work (Cunningham et al., 2024; Templeton et al., 2024).
@@ 3 How does specialization behave?
We begin in this section by developing a mechanistic understanding of how TTT behaves. Since we do not have access to the “true” hypothesized concept space $\Phi$, we train SAEs to learn a concept space $\smash{\hat{\Phi}}$ whose characteristics can be probed. We employ a top-$k$ SAE (Gao et al., 2025) to learn sparse feature representations. After a brief description of the experimental setup, we detail our key observations O1-O3 from Section˜1.
@@ 3.1 Experimental setup
We use the ImageNet-1K dataset (Deng et al., 2009). The dense vectors $\Psi(x)$ are normalized CLIP embeddings (Radford et al., 2021) of dimension $d_{2}=512$ (we use only the <CLS> component). We set the sparse dimension to $d_{1}=8\times d_{2}=4096$ and the sparsity level to $s=16$. For our analysis, we use the SAE’s reconstructions $\smash{\hat{\Psi}(x)}$ rather than the original CLIP embeddings $\Psi(x)$. This choice aligns our experiments more closely with our theoretical model and circumvents known challenges in training SAEs on raw, complex embeddings. This comes at the cost of a mild $6\%$ drop in accuracy for a global linear classifier trained on the embeddings. [^3] Note that the SAE is trained in an unsupervised way, without explicitly retaining classification accuracy, yet using the SAE’s features leads only to a minor drop in accuracy. In Section˜E.2, we additionally validate the SAE framework on MNIST (LeCun et al., 1998b), where we use $\Psi(x)$ directly rather than its reconstruction, as recovering an accurate $\smash{\hat{\Phi}(x)}$ is much easier on this simpler dataset. We include training details in Section˜E.3.
@@ 3.2 Results
With a sparsity penalty of $\lambda=0.2$, the learned masks are highly sparse, activating on average only $\|m\|_{0}\approx 40$ concepts. This is substantially smaller than the total number of unique concepts active across the neighborhood, which is approximately $180$. As shown in Table˜1 (TTT column), this sparsely supported model performs on par with TTT on top of dense reconstructions $\smash{\hat{\Psi}(x)}$ [^5] TTT in the concept space without extra masking achieves similar results. The fact that it also works with masking provides strong evidence that TTT adjusts only a few selected concepts in the neighborhood.. This suggests that a small, adaptively chosen set of concepts is sufficient to capture the relevant information within a local region. We obtain similar results for the Gemma Scope SAE (Lieberum et al., 2024) on MNIST data, which we present in Section˜E.2.
@@
Figure 3: Comparison of predicted class probabilities for TTT models trained on dense reconstructions ($\smash{\hat{\Psi}}$) and sparse concepts ($\smash{\hat{\Phi}}$), relative to their magnitude. The small relative TV distance ($\ll\!\!1$; defined in Section˜E.1) between both distributions indicates strong functional agreement between TTT in $\smash{\hat{\Psi}}$ and $\smash{\hat{\Phi}}$ space. We show 90% bootstrap confidence intervals across $1000$ test points.
@@
While the adaptive masking in Equation˜3 explicitly enforces a sparse solution, we find evidence that standard TTT in the feature space implicitly favors a solution that is sparse in the concept space. First, the TTT models trained on dense reconstructions $\smash{\hat{\Psi}(x)}$ and sparse concepts $\smash{\hat{\Phi}_{m}(x)}$ achieve nearly identical accuracy (cf. Table 1). Furthermore, their predictions agree in $\approx\!\!89\%$ of cases, indicating that they learn functionally equivalent classifiers (apart from pathological examples). Figure˜3 reinforces this by showing that both models lead to closely matched predictive distributions over the top-10 predicted classes. In Figure˜3, we compare the ordered predicted probabilities for $\smash{\hat{\Phi}}$ to the corresponding probabilities for $\smash{\hat{\Psi}}$, matching the distributions’ temperatures, and averaging over all test points. Their strong correspondence suggests that TTT on reconstructed embeddings is implicitly biased towards a sparse solution in the underlying concept space. This phenomenon may be linked to the implicit bias of optimization algorithms (e.g., SGD or Adam), which are known to favor minimum-norm solutions (Gunasekar et al., 2018; belkin2019reconciling; frei2022benign). When the feature map superimposes concepts, this implicit bias may favor sparse solutions in the underlying concept space (Vaskevicius et al., 2019).
@@ 4.1 Settings
We consider the following three tasks, and provide implementation details in Appendix˜E:
  • MNIST. We train a global classifier following the LeNet-5 (LeCun et al., 1998a) architecture with a cross-entropy loss on the MNIST (LeCun et al., 1998b) dataset of handwritten digits. To distinguish learning of concepts from TTT’s ability to identify the correct weights for each locally relevant concept, we train only the last linear layer during TTT. The last linear layer is updated to minimize the cross-entropy loss on the neighborhood of a test point (cf. Equation˜2). We report classification error on the test set.
@@ 4.2 Results
We perform a scaling study by varying the model size and comparing the performance of TTT against global training and a majority vote on the neighborhood (i.e., a simple non-parametric baseline). Figure˜4 shows the results. In MNIST, we train convolutional neural networks of varying sizes as summarized in Appendix˜E. In ImageNet, we train multi-layer perceptrons on top of CLIP embeddings, varying the hidden dimension. In language modeling, we evaluate Qwen2.5 base models of sizes ranging from 0.5B to 32B parameters. We find across all tasks that TTT outperforms global training and majority vote, with the performance gap shrinking as the model size increases. We hypothesize that at a larger model size, fewer concepts have to be superimposed in latent space, leading to less interference when globally mapping latent representations to predictions. While a larger model size allows for better global disentanglement of concepts, TTT can compensate for limited model capacity by adapting the head to the specific concepts in a local neighborhood.
@@
Next to performing a scaling study on model size, we also vary the dataset size. Figure˜5 shows the results at a fixed model scale. We subsample the training datasets of ImageNet and MNIST to fractions ranging from $1\%$ to $100\%$ of the original training set size, ensuring that all subsampled datasets are class-balanced. We then train global models and evaluate TTT on the respective test sets. We find that TTT consistently outperforms global training, with the slight trend of the performance gap widening as the dataset size increases. We hypothesize that larger datasets provide richer neighborhoods for local adaptation, enabling TTT to specialize more effectively to the specific concepts relevant to each test point.
@@ 5.1 Analyzing the in-distribution test error of TTT
In Section˜3, we have made the surprising observation that the neighborhood of a test point $x^{\star}$ is explained by only a few active concepts (cf. O2). Based on this, we assume that locally there exists an $\Theta(s)$-sparse concept vector $w_{x^{*}}\in\mathbb{R}^{d_{1}}$ that approximates $\langle\Phi(x),w_{\star}\rangle$ over the test point’s neighborhood. This assumption may seem unrealistic at first, since the neighborhood of $x^{\star}$ has $k$ samples and up to $s\cdot k$ active concepts. Yet, as we observe empirically in Section˜3, TTT learns a $\Theta(s)$-sparse regressor that does well on the entire neighborhood.
@@
With these hypotheses, we can bound the in-distribution test error of TTT using techniques from sparse recovery (Bickel et al., 2009; Van De Geer & Bühlmann, 2009):
@@
Let $\Phi:\mathcal{X}\to\mathbb{R}^{d_{1}}$ be an $s$-sparse concept space and $\Psi:\mathcal{X}\to\mathbb{R}^{d_{2}}$ be a learned feature map with $d_{2}\ll d_{1}$. Let $f(x)=\langle\Phi(x),w_{\star}\rangle$ be the ground truth function and data be $\sigma^{2}$-subgaussian. Let $x^{\star}$ be a test point with a neighborhood (in $\Psi$-space) of sufficiently small size $k$ such that Hypotheses 1–3 hold. Let the learned features $\Psi$ be sufficiently expressive to represent $f$ locally (in particular, $d_{2}\geq\Omega(s\log d_{1})$). We denote by $\hat{v}_{x^{\star}}^{\text{TTT}}$ the local empirical risk minimizer on the neighborhood of $x^{\star}$.
@@
We expand on our theoretical results in the appendix as follows: In addition to increasing the effective model capacity, we also explore whether one can understand TTT under the LRH through the lens of statistical learning theory. Specifically, in Section˜A.1, we explore this direction using notions from low-degree polynomials and hypercontractivity (Klivans et al., 2008; Paouris et al., 2022; Damian et al., 2024; Bizeul & Klartag, 2025, and references therein). Moreover, in Section˜A.2, we contrast TTT to classical non-parametric methods (Fix & Hodges Jr., 1951; Nadaraya, 1964; Watson, 1964) such as majority voting, which underperform in our experiments (cf. Section˜4).

### v3

@@ 1 Introduction
Since the “ImageNet moment” in 2012 when AlexNet won the ImageNet challenge (Krizhevsky et al., 2012), scaling data, parameters, and compute have led to foundation models, which are large-scale neural networks trained on vast datasets that can be applied across a wide range of use cases (Bommasani et al., 2021). This has spurred research on scaling laws, suggesting that scaling pre-training of a single model on a broad data distribution is sufficient for good performance on downstream tasks (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022). With first-generation foundation models, fine-tuning was used primarily to adapt models to out-of-distribution test data (i.e., with a distribution shift) or to leverage fresh training data that was not seen during pre-training (so-called “privileged” data). Test-time training (TTT; Sun et al., 2020; Hardt & Sun, 2024; Akyürek et al., 2025) emerged as pushing this mechanism to the extreme: fine-tuning a separate model for each prediction. In recent years, foundation models have grown so large that most test data is effectively “in-distribution”, meaning the model has encountered similar data during pre-training. This raises a key question:
@@
While foundation models typically have a large parameter count, extensive results on scaling laws show continuing improvements in performance when scaling model size (Kaplan et al., 2020; Bubeck & Sellke, 2021). Consequently, our work posits that today’s foundation models are effectively “underparameterized”, meaning that performance can be improved by further scaling. We hypothesize that due to this effective underparameterization, even if test data is in-distribution, the model cannot simultaneously approximate the ground truth across the full data distribution. TTT offers a mechanism to specialize the model to a local area around the test example. By temporarily “forgetting” irrelevant pre-trained knowledge, the model “frees up” capacity to learn the relevant concepts to the immediate task at a higher resolution. We refer to this mechanism as specialization after generalization. The mechanism of TTT—temporarily reallocating capacity by “forgetting” irrelevant knowledge—connects to concepts of capacity saturation and interference studied in continual learning (McCloskey & Cohen, 1989; Kirkpatrick et al., 2017).
@@
The LRH can be formalized as follows. We assume the existence of an $s$-sparse concept space $\Phi:\mathcal{X}\to\mathbb{R}^{d_{1}}$} which is approximated by a learned feature map $\Psi:\mathcal{X}\to\mathbb{R}^{d_{2}}$} with $d_{2}\ll d_{1}$. The ground truth to be learned is linear in the concept space, i.e., $f^{*}(x)=\langle\Phi(x),w_{\star}\rangle$ for some unknown $w_{\star}\in\mathbb{R}^{d_{1}}$. [^1] In classification, the logits are canonically parameterized as $\langle\Phi(x),w_{c}\rangle$ where $w_{c}$ is the class-specific weight vector. The class probability is then given by $\mathbb{P}(c\mid x)\propto\exp(\langle\Phi(x),w_{c}\rangle)$. In Section˜3, we leverage the LRH to develop a mechanistic understanding of how TTT behaves, and make the following key observations:
@@ 2. O2:
    Among a test point $x^{\star}\in\mathcal{X}$ and its neighborhood (in $\Psi$-space), $f^{*}$ can be approximated by an $s$-sparse linear function in the concept space $\Phi$.
@@ 2 Related work
In the classical machine learning paradigm, models are trained on a fixed training set and then kept frozen during evaluation. Despite this standard practice that was used for decades, early work suggested specializing the model at test-time to each prediction task—such examples are local learning (Cleveland, 1979; Cleveland & Devlin, 1988; Atkeson et al., 1997) and local fine-tuning (Bottou & Vapnik, 1992). More recently, the idea of TTT (Sun et al., 2020; Wang et al., 2021) has regained attention in the context of fine-tuning large foundation models during evaluation (e.g., Krause et al., 2018; Hardt & Sun, 2024; Sun et al., 2025). TTT for a few gradient steps on (self-)supervised losses has since shown success in domains such as control (Hansen et al., 2021), abstract reasoning (Akyürek et al., 2025; Zweiger et al., 2025), language modeling (Hardt & Sun, 2024; Hübotter et al., 2025; Sun et al., 2025; Zhang et al., 2025; von Oswald et al., 2025; Yu et al., 2025), and video generation (Dalal et al., 2025). Many standard TTT methods train on carefully selected data from the pre-training dataset (i.e., do not add any new privileged information; Hardt & Sun, 2024; Hübotter et al., 2025), and several works studied how to optimally select data for imitation, e.g., the early seminal work of MacKay (1992) and recent extensions (Hübotter et al., 2024; Bagatella et al., 2025b). TTT has also been extended from supervised learning to reinforcement learning (Zuo et al., 2025; Bagatella et al., 2025a; Diaz-Bone et al., 2025).
@@
Our theoretical framework is built upon the LRH, which posits that foundation models represent high-level concepts as linear directions in their activation spaces. This idea has its roots in early word embedding models, which famously showed that semantic analogies could be solved with simple vector arithmetic (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016). More recently, the LRH has been validated across a wide range of models and domains, with studies identifying linear representations for abstract concepts like sentiment (Tigges et al., 2023), the state of a game board (Nanda et al., 2023), and even fundamental axes of space and time (Gurnee & Tegmark, 2024). A key tool for discovering and studying these conceptual directions is SAEs (Makhzani & Frey, 2014; Lieberum et al., 2024; Gao et al., 2025). SAEs are auxiliary models trained to reconstruct a foundation model’s internal activations from a sparse, overcomplete dictionary of features. This process often yields features that are monosemantic, or aligned with single, human-interpretable concepts, thereby providing an empirical method to uncover the sparse concept space we consider in our work (Cunningham et al., 2024; Templeton et al., 2024).
@@ 3 How does specialization behave?
In this section, we begin by developing a mechanistic understanding of how TTT behaves. Since the “true” hypothesized concept space $\Phi$ is not accessible, we train SAEs to learn an approximate concept space $\smash{\hat{\Phi}}$ whose properties can be analyzed. We use a top-$k$ SAE (Gao et al., 2025) to obtain sparse feature representations. After a brief description of the experimental setup, we detail our key observations O1-O3 from Section˜1.
@@ 3.1 Experimental setup
We use the ImageNet-1K dataset (Deng et al., 2009). The dense vectors $\Psi(x)$ are normalized CLIP embeddings (Radford et al., 2021) of dimension $d_{2}=512$ (we use only the <CLS> component). We set the sparse dimension to $d_{1}=8\times d_{2}=4096$ and the sparsity level to $s=16$. For our analysis, we use the SAE’s reconstructions $\smash{\hat{\Psi}(x)}$ rather than the original CLIP embeddings $\Psi(x)$. This choice aligns our experiments more closely with our theoretical model and circumvents known challenges in training SAEs on raw, complex embeddings. This comes at the cost of a mild $6\,\%$ drop in accuracy for a global linear classifier trained on the embeddings. [^2] Note that the SAE is trained in an unsupervised way, without explicitly retaining classification accuracy, yet using the SAE’s features leads only to a minor drop in accuracy. In Section˜F.2, we additionally validate the SAE framework on MNIST (LeCun et al., 1998b), where we use $\Psi(x)$ directly rather than its reconstruction, as recovering an accurate $\smash{\hat{\Phi}(x)}$ is much easier on this simpler dataset. We include training details in Section˜F.3.
@@ 3.2 Results
With a sparsity penalty of $\lambda=0.2$, the learned masks are highly sparse, activating on average only $\|m\|_{0}\approx 40$ concepts. This is substantially smaller than the total number of unique concepts active across the neighborhood, which is approximately $180$. As shown in Table˜1 (TTT column), this sparsely supported model performs on par with TTT on top of dense reconstructions $\smash{\hat{\Psi}(x)}$ [^4] TTT in the concept space without extra masking achieves similar results. The fact that it also works with masking provides strong evidence that TTT adjusts only a few selected concepts in the neighborhood.. This suggests that a small, adaptively chosen set of concepts is sufficient to capture the relevant information within a local region. We obtain similar results for the Gemma Scope SAE (Lieberum et al., 2024) on MNIST data, which we present in Section˜F.2.
@@
Figure 3: Comparison of predicted class probabilities for TTT models trained on dense reconstructions ($\smash{\hat{\Psi}}$) and sparse concepts ($\smash{\hat{\Phi}}$), relative to their magnitude. The small relative TV distance ($\ll\!\!1$; defined in Section˜F.1) between both distributions indicates strong functional agreement between TTT in $\smash{\hat{\Psi}}$ and $\smash{\hat{\Phi}}$ space. We show 90 % bootstrap confidence intervals across $1000$ test points.
@@
While the adaptive masking in Equation˜2 explicitly enforces a sparse solution, we find evidence that standard TTT in the feature space implicitly favors a solution that is sparse in the concept space. First, the TTT models trained on dense reconstructions $\smash{\hat{\Psi}(x)}$ and sparse concepts $\smash{\hat{\Phi}_{m}(x)}$ achieve nearly identical accuracy (cf. Table 1). Furthermore, their predictions agree in $\approx\!\!89\,\%$ of cases, indicating that they learn functionally equivalent classifiers (apart from pathological examples). Figure˜3 reinforces this by showing that both models lead to closely matched predictive distributions over the top-10 predicted classes. In Figure˜3, we compare the ordered predicted probabilities for $\smash{\hat{\Phi}}$ to the corresponding probabilities for $\smash{\hat{\Psi}}$, matching the distributions’ temperatures, and averaging over all test points. Their strong correspondence suggests that TTT on reconstructed embeddings is implicitly biased towards a sparse solution in the underlying concept space. This phenomenon may be linked to the implicit bias of optimization algorithms (e.g., SGD or Adam), which are known to favor minimum-norm solutions (Gunasekar et al., 2018; Belkin et al., 2019; Frei et al., 2022). When the feature map superimposes concepts, this implicit bias may favor sparse solutions in the underlying concept space (Vaskevicius et al., 2019).
@@ 4.1 Settings
We consider the following three tasks, and provide implementation details in Appendix˜F:
  • MNIST. We train a global classifier following the LeNet-5 (LeCun et al., 1998a) architecture with a cross-entropy loss on the MNIST (LeCun et al., 1998b) dataset of handwritten digits. To distinguish learning of concepts from TTT’s ability to identify the correct weights for each locally relevant concept, we restrict TTT to updating only the last linear layer. This layer is fine-tuned to minimize the cross-entropy loss on the neighborhood of each test point (cf. Equation˜1), while all earlier layers remain fixed. As shown by additional experiments in Section˜D.2, end-to-end TTT does not yield a meaningful performance improvement, validating this design choice. We report classification error on the test set.
@@ 4.2 Results
We conduct a scaling study by varying the model size and comparing the performance of TTT against global training as well as a majority vote baseline over the neighborhood (i.e., a simple non-parametric approach). The results are shown in Figure˜4. For MNIST, we train convolutional neural networks of varying sizes, as summarized in Appendix˜F. For ImageNet, we train multi-layer perceptrons on top of CLIP embeddings, varying the hidden dimension. In language modeling, we evaluate Qwen2.5 base models of sizes ranging from 0.5B to 32B parameters. We find across all tasks that TTT outperforms global training and majority vote, with the performance gap shrinking as the model size increases. We hypothesize that at a larger model size, fewer concepts have to be superimposed in latent space, leading to less interference when globally mapping latent representations to predictions. While a larger model size allows for better global disentanglement of concepts, TTT can compensate for limited model capacity by adapting the head to the specific concepts in a local neighborhood.
@@
Next to performing a scaling study on model size, we also vary the dataset size. Figure˜5 shows the results at a fixed model scale. We subsample the training datasets of ImageNet and MNIST to fractions ranging from 11\, to $100\,\%$ of the original training set size, ensuring that all subsampled datasets are class-balanced. We then train global models and evaluate TTT on the respective test sets. We find that TTT consistently outperforms global training, with the slight trend of the performance gap widening as the dataset size increases. We hypothesize that larger datasets provide richer neighborhoods for local adaptation, enabling TTT to specialize more effectively to the specific concepts relevant to each test point.
@@ 5.1 Analyzing the in-distribution test error of TTT
In Section˜3, we have made the surprising observation that the neighborhood of a test point $x^{\star}$ is explained by only a few active concepts (cf. O2). Based on this, we assume that locally there exists an $\Theta(s)$-sparse concept vector $w_{x^{*}}\in\mathbb{R}^{d_{1}}$ that approximates $\langle\Phi(x),w_{\star}\rangle$ over the test point’s neighborhood. This assumption may seem unrealistic at first, since the neighborhood of $x^{\star}$ has $k$ samples and up to $s\cdot k$ active concepts. Yet, as we observe empirically in Section˜3, TTT learns a $\Theta(s)$-sparse regressor that does well on the entire neighborhood if the neighborhood size is sufficiently small.
@@
With these hypotheses, we can bound the in-distribution test error of TTT using techniques from sparse recovery (Bickel et al., 2009; Van De Geer & Bühlmann, 2009): [^7] Our results indicate that TTT may not converge quickly (or at all) if the label noise $\sigma^{2}$ is large or any of Hypotheses 1-3 were to be violated. Notably, the neighborhood size $k$ needs to be sufficiently small for Hypothesis 2 to hold.
@@
Let $\Phi:\mathcal{X}\to\mathbb{R}^{d_{1}}$ be an $s$-sparse concept space and $\Psi:\mathcal{X}\to\mathbb{R}^{d_{2}}$ be a learned feature map with $d_{2}\ll d_{1}$. Let $f(x)=\langle\Phi(x),w_{\star}\rangle$ be the ground truth function and labels be $\sigma^{2}$-subgaussian. Let $x^{\star}$ be a test point with a neighborhood (in $\Psi$-space) of sufficiently small size $k$ such that Hypotheses 1–3 hold. Let the learned features $\Psi$ be sufficiently expressive to represent $f$ locally (in particular, $d_{2}\geq\Omega(s\log d_{1})$). We denote by $\hat{v}_{x^{\star}}^{\text{TTT}}$ the local empirical risk minimizer on the neighborhood of $x^{\star}$.
@@
We expand on our theoretical results in the appendix as follows: First, we explore whether one can understand TTT under the LRH through the lens of statistical learning theory. Specifically, in Section˜A.1, we explore this direction using notions from low-degree polynomials and hypercontractivity (Klivans et al., 2008; Paouris et al., 2022; Damian et al., 2024; Bizeul & Klartag, 2025, and references therein). Additionally, in Section˜A.2, we contrast TTT to classical non-parametric methods (Fix & Hodges Jr., 1951; Nadaraya, 1964; Watson, 1964) such as majority voting, which underperform in our experiments (cf. Section˜4).